11-You have decided to buy a used car. The dealer has offered you two options: (FV of $1, PV of $1, FVA of $1, and PVA of $1) (Use the appropriate factor(s) from the tables provided.) Pay $650 per month for 20 months and an additional $12,000 at the end of 20 months. The dealer is charging an annual interest rate of 24%. Make a one-time payment of $17,354, due when you purchase the car. 1-a. Determine how much cash the dealer would charge in option (a). (Round your answer to 2 decimal places.) 1-b. In present value terms, which offer is clearly a better deal?
Answer #1

1-a Total cash paid = $650*20 + 12,000 = $25,000

Monthly interest rate = 24%/12 = 2%

Present value = 650*PVAF(2%, 20 periods) + 12,000*PVF(2%, 20 periods)

= 650*16.351 + 12,000*0.673

= $18,704.15

One time payment is a better deal